An investor holds shares i three independents companies, companies A, B and C. An analyst predicts the probability of profit increases for each company as follows:
Company A: 0.70
Company B: 0.50
Company C: 0.65
1. If the analyst's predictions are correct, what is the probability that all 3 companies show a profit increase? (express your answer to the nearest hundredth of one percent)
2. If the analyst's predictions are true, what is the probability that at least one company shows a profit increase?(express your answer to the nearest hundredth of one percent)

Answer by stanbon(75887)

An investor holds shares i three independents companies, companies A, B and C. An analyst predicts the probability of profit increases for each company as follows:
Company A: 0.70
Company B: 0.50
Company C: 0.65
1. If the analyst's predictions are correct, what is the probability that all 3 companies show a profit increase? (express your answer to the nearest hundredth of one percent)
Answer: P(all are profitable)= 0.7*0.5*0.65=0.2275=22.75%
-----------------------
2. If the analyst's predictions are true, what is the probability that at least one company shows a profit increase?(express your answer to the nearest hundredth of one percent)
Answer: Prob (at least one profits)= 1-P(all lose money)
=1-(0.3)().5)(0.35)= 0.9475=94.75%
----------------
Cheers,
Stan H.